# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v3

@@ Abstract
Because it is difficult to precisely specify complex objectives, reinforcement learning policies are often optimized using proxy reward functions that only approximate the true goal. However, optimizing proxy rewards frequently leads to reward hacking: the optimized reward function ceases to be a good proxy and the resulting policy performs poorly with respect to the unspecified true reward. Principled solutions to reward hacking have been impeded by the lack of a good definition for the problem. To address this gap, we introduce a definition of reward hacking based on the correlation between proxy and true rewards for states and actions seen by a “base policy” that breaks down under optimization. We show that this definition captures reward hacking behavior across several realistic settings, including in reinforcement learning from human feedback (RLHF). Using our formulation, we show theoretically that regularization to the base policy can effectively prevent reward hacking. While the current practice in RLHF applies a KL penalty between action distributions for this purpose, our theory suggests regularizing the $\chi^{2}$ divergence between the policies’ occupancy measures can be more effective. We intuitively show the benefits of this type of regularization and demonstrate that it better mitigates reward hacking in practice across four realistic settings, including RLHF. Our code is available at https://github.com/cassidylaidlaw/orpo.
@@ 1 Introduction
Since proxy reward functions for complex tasks are nearly always misspecified in practice, what can be done to avoid reward hacking? There is a lack of principled solutions for preventing reward hacking, stemming from a more fundamental problem: defining reward hacking in a formal sense that captures realistic cases. In particular, it is difficult to characterize what makes a proxy reward function “good”. For example, Skalse et al. (2022) define a notion of “unhackability” but find that it can only hold if such an objective is a multiple of the true reward—a requirement that is unlikely in practice.
Figure 1: We present a new characterization of reward hacking and a method for preventing it. We define a proxy reward function as one that correlates with an unknown true reward function for state-action pairs sampled from some base policy. However, optimizing the proxy alone can lead to a breakdown in the correlation and worse true reward than the base policy. We show theoretically and empirically that optimizing the proxy with $\chi^{2}$ occupancy measure regularization to the base policy can allow outperforming the base policy under the unknown true reward.
@@
We argue that proxies are chosen because they seem to capture the true objective under some base distribution of behavior. To formalize this, we define a proxy reward as one that correlates with the true reward function under the distribution of states and actions visited by a base policy. Then, we define a hackable proxy as one that induces a decrease in true reward compared to the base policy when optimized (Figure 1). This framework captures a wide range of intuitive reward hacking cases across realistic environments: traffic control, pandemic mitigation, blood glucose regulation, and reinforcement learning from human feedback (RLHF) (Figure 2).
Furthermore, our definition motivates a method for avoiding reward hacking during policy optimization by regularizing the optimized policy to be similar to the base policy. Specifically, we find that optimizing the proxy reward minus a regularization term provides a provable lower bound on improvement in true reward. The amount of regularization needed depends on the correlation between the proxy and true rewards; as the correlation becomes stronger, it is possible to use less regularization. In practice, RLHF already uses this form of regularization through a KL divergence penalty to the base policy (Stiennon et al., 2020; Bai et al., 2022), and Theorem 5.1 provides theoretical justification for why this works. However, our results suggest an important modification: rather than penalizing the KL divergence between the policies’ action distributions as currently done in RLHF, we should instead penalize the chi-squared ($\chi^{2}$) divergence between the policies’ occupancy measures. A policy’s occupancy measure (OM) is the distribution of states or state-action pairs seen by the policy when it interacts with its environment. Unlike action distribution (AD)-based metrics, occupancy measures capture not just the actions an agent takes, but also the states it reaches. We compare both OM to AD regularization and $\chi^{2}$ to KL divergence, presenting intuitive reasons why $\chi^{2}$ OM divergence may be a better regularization target in Figures 3 and 4.
Based on our theoretical results, we empirically investigate the benefits of using $\chi^{2}$ occupancy measure regularization to prevent reward hacking. Following previous work (Ho & Ermon, 2016), we implement OM regularization using a discriminator network that approximates the OM divergence between policies. We evaluate our approach in multiple reward hacking benchmark environments (Pan et al., 2022), comparing the effects of OM and AD regularization with both $\chi^{2}$ and KL divergence on policies optimized with hackable proxy reward functions. The results of our experiments demonstrate that training with occupancy measure regularization leads to better performance under the unseen true reward function in all of the environments, validating our theoretical results. In contrast, we find that it is difficult to tune AD regularization in some environments to both prevent reward hacking and allow meaningful improvement over the base policy. Furthermore, regularization with $\chi^{2}$ divergence tends to be more robust to the choice of regularization coefficient and often achieves higher true reward than regularizing with KL divergence.
@@
    We provide a new formal definition of reward hacking and show that it captures a number of realistic case studies.
@@
    Using our definition, we establish that optimizing a proxy reward with $\chi^{2}$ occupancy measure (OM) regularization leads to a provable improvement in the unknown true reward function.
@@ 2 Related Work
While there have been separate lines of work investigating reward hacking and exploring the use of occupancy measure divergences for other applications, to the best of our knowledge, we are the first to specifically study the application of occupancy measure regularization to the problem of reward hacking.
Reward hacking.  Some prior works establish theoretical models of reward hacking as a special case of Goodhart’s Law (Goodhart, 1984; Leike et al., 2018; Manheim & Garrabrant, 2019; Krakovna, 2019; Skalse et al., 2022; Ngo et al., 2023). Krakovna (2018) catalog many examples of reward hacking. Pan et al. (2022) categorize different types of reward misspecification and relate optimization power to reward hacking.
Safe reinforcement learning.  In the context of reward hacking, regularizing policies to be similar to an offline policy based on their action distribution KL divergence was proposed by Stiennon et al. (2020) and has since been widely employed in the context of optimizing LLMs using RLHF (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022). KL regularization for RLHF has been further studied by Vieillard et al. (2021), Gao et al. (2022), and Korbak et al. (2022). Nika et al. (2024) propose a type of occupancy measure regularization in RLHF but it is limited to deterministic MDPs, while we study general stochastic MDPs. Some alternative approaches to avoid reward hacking include quantilizers (Taylor, 2016), “mild” optimization (Taylor et al., 2020), and impact regularization (Turner et al., 2020). While constrained RL can prevent the misbehavior of agents that optimize flawed reward functions (Dalal et al., 2018; Chow et al., 2019; Zhang et al., 2020; Roy et al., 2022), it simply shifts the difficulty of designing a reward function to specifying a set of constraints and weights. Robust RL usually considers a misspecified transition model, but some work has explored misspecified reward functions (Derman et al., 2021; Gadot et al., 2024). Other proposals to address the reward specification problem attempt to infer the true reward function based on the given proxy reward function, environment context, and/or feedback from humans (Hadfield-Menell et al., 2017; Reddy et al., 2020; Lee et al., 2021). Gleave et al. (2021) have previously studied quantifying the similarity of reward functions.
@@
Applications of occupancy measure regularization.  Occupancy measure regularization and optimization have been used for a variety of purposes. GAIL (Ho & Ermon, 2016) is an algorithm for robust imitation learning that aims to match the imitator’s occupancy measure to that of the demonstrator. Kang et al. (2018) combines GAIL with a reward function to efficiently explore using human data. Another line of work aims to find a policy with the highest-entropy occupancy measure for the purpose of exploring the state space (Hazan et al., 2019; Lee et al., 2020; Nedergaard & Cook, 2023). Various types of distributional regularization are used in model-based RL since learned models may not generalize out-of-distribution (Yang et al., 2022).
@@
Refer to caption04812-100-10-50Basepolicy(4.5, -7.6)Trafficcontrol(r=0.10)Refer to caption04812-100-10-50Rewardhacking policyTrue reward $R(s,a)$(9.1, -197.2)Refer to caption04812-100-10-50Regularizedpolicy(4.9, -4.2)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.06, -0.06)Pandemicmitigation(r=0.75)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.01, -0.14)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.05, -0.05)Refer to caption-0.2-0.10-1000-100-500(-0.04, -28)Glucosemonitoring(r=0.70)Refer to caption-0.2-0.10-1000-100-500(-0.00, -106)Refer to caption-0.2-0.10Proxy reward $\tilde{R}(s,a)$-1000-100-500(-0.01, -8)Refer to caption-3030102030(-0.0, 16.4)RLHF(r=0.40)Refer to caption-3030102030(1.4, 10.0)Refer to caption-3030102030(0.3, 17.0)Refer to caption00.20.400.050.1(0.07, 0.06)AI safetygridworld(r=0.40)Refer to caption00.20.400.050.1(0.45, 0.02)Refer to caption00.20.400.050.1(0.09, 0.09)Mean reward
Figure 2: Our definition of reward hacking successfully describes reward hacking behavior in four realistic environments and an illustrative gridworld. The top row shows the distribution of proxy and true reward values for state-action pairs sampled from a domain-appropriate base policy for each environment; in all environments, the proxy and true rewards are correlated. However, as shown in the middle row, this correlation breaks down if the proxy is optimized via RL, and the true reward achieved is lower than that of the base policy, which we characterize as reward hacking. In line with our theoretical results, the bottom row shows that RL with occupancy measure regularization to the base policy can prevent reward hacking while enabling an increase in true reward.
@@ 3 Preliminaries
Intuitively, the occupancy measure represents the distribution of states and actions visited by the policy over time. Furthermore, combining (1) and (2), we can see that the return of a policy can be re-written as the expected reward for states and actions sampled from the occupancy measure:
@@
Prior definitions of reward hacking.  Due to these challenges, prior work has struggled to define reward hacking. Skalse et al. (2022) attempted to address our first desideratum of defining reasonable proxies by introducing a notion of an “unhackable” proxy reward. However, they found that a proxy is only unhackable if it is equivalent to the true reward up to scaling—a definition that is clearly too restrictive. The robust MDP literature has considered proxy rewards that differ from the true reward by at most some constant, e.g., $|R(s,a)-\tilde{R}(s,a)|\leq c$ for all $s,a$ (Derman et al., 2021). However, we find that in many realistic cases of reward hacking, proxies can differ from the true reward by arbitrarily large amounts in certain states, making these definitions inapplicable. As for our second desideratum of defining a true reward threshold, there is little existing literature.
@@ 4 Defining Reward Hacking
Despite the difficulties in formalizing reward hacking, we argue that both of our desiderata for a definition can be met by defining reward hacking with respect to a base policy. This approach enables both a precise definition of proxy rewards and a natural true reward threshold for determining when reward hacking is occurring.
Characterizing proxy rewards.  To find a realistic definition of a good proxy reward, consider the process by which system designers create proxies. If they are hand-specified, we argue that designers usually imagine a distribution of behavior and design a reward that captures the objective under that distribution. For example, we study a traffic control simulator (Lopez et al., 2018; Vinitsky et al., 2018; Wu et al., 2022; Pan et al., 2022) where a small number of autonomous cars help to regulate traffic among a larger population of human drivers. In this case, a designer might choose average vehicle speed as a proxy for traffic flow; this is a reasonable proxy because under the distribution of typical human driving behavior, higher speeds are associated with better traffic conditions.
We can formalize this intuition by requiring that the proxy and true rewards be correlated over states and actions sampled from a base policy:
@@
An $r$-correlated proxy reward $\tilde{R}$ with respect to a base policy $\pi_{\text{base}}$} is one that has a correlation of $r>0$ with the true reward for state-action pairs sampled from the base policy:
@@
$\textstyle\qquad\qquad\quad\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\left(\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}\right)\left(\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\right)\right]=r,$
@@
$\textstyle\text{where}\quad\sigma_{\tilde{R}}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})\big)^{2}\right]\quad\text{and}\quad\sigma_{R}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(R(s,a)-J({\pi_{\text{base}}},R)\big)^{2}\right]$
are the variances of proxy and true rewards under the base policy.
@@
To consider an application of our definition, let’s return to the example of our traffic environment, where we define the true reward function as the negative total commute time for all vehicles. Using an autonomous vehicle policy based on a common model of human driving behavior as $\pi_{\text{base}}$}, we plot the distribution of true and proxy rewards in the top-left corner of Figure 2. The clear correlation that we observe between the two reward functions—higher average speed corresponding to lower commute times and vice versa—confirms that average speed is a correlated proxy reward according to Definition 4.1. This demonstrates that our formal definition successfully captures intuitive real-world proxy rewards.
@@
Definition 4.1 offers key advantages over past formalisms. First, it avoids too strongly constraining proxy rewards by allowing the proxy and true reward functions to diverge arbitrarily at some state-action pairs when those pairs have low or zero occupancy measure under the base policy. Second, it encompasses both learned proxy rewards and hand-specified ones. In reward learning, we collect annotations (such as preference comparisons or scalar feedback) for states and actions collected from a rollout policy (such as the SFT model in RLHF). Then, a reward function is estimated via supervised learning over this dataset with an appropriate loss function. If the estimated reward function generalizes well in-distribution, then we should expect the true and learned rewards to be well-correlated under states and actions sampled from the rollout policy, satisfying Definition 4.1 (see Lemma A.8).
@@
Choosing a threshold for reward hacking.  As discussed in Section 3, defining reward hacking requires specifying a threshold of true reward below which policy behavior is considered poor enough to be classified as reward hacking. Given that we already characterize proxy rewards with respect to a base policy, this same base policy provides a natural baseline for evaluating a policy that optimizes the proxy: if optimizing the proxy leads to worse true reward than the base policy achieves, then the system designer may be better off simply using the base policy.
Definition 4.2 (Hackable proxy reward).
@@
Suppose $\tilde{R}$ is an $r$-correlated proxy with respect to $\pi_{\text{base}}$} (Definition 4.1). Then we say reward hacking occurs when a policy $\pi$ optimized for $\tilde{R}$ has lower true reward than the base policy $\pi_{\text{base}}$}, i.e., when $J(\pi,R)<J({\pi_{\text{base}}},R)$. If an optimal policy for $\tilde{R}$ exhibits reward hacking, then we say that the proxy reward $\tilde{R}$ is hackable:
@@
For example, in the traffic control environment, optimizing average speed as a proxy reward leads to autonomous vehicles blocking highway on-ramps, a strategy that increases overall average speed by allowing highway traffic to move freely while forcing on-ramp cars to stop completely. However, the true reward becomes extremely low because commute times for the cars on the on-ramp are now arbitrarily long. This policy is thus reward hacking under Definition 4.2 since the true reward that it accumulates is lower than that achieved by typical human driving.
@@
Verifying our definition experimentally.  To test whether Definition 4.2 accurately captures intuitive cases of reward hacking in practice, we consider four realistic environments and an illustrative gridworld. In addition to the aforementioned traffic simulator, we consider a pandemic mitigation and glucose monitoring environments, all of which were originally studied by Pan et al. (2022) as examples of reward hacking. PandemicSimulator (Kompella et al., 2020) uses a specialized SEIR (susceptible, exposed, infected, recovered) infection model to simulate the COVID-19 pandemic among a population; the policy controls the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward omits the political cost of certain decisions. SimGlucose (Man et al., 2014) is based on an FDA-approved simulator of Type 1 Diabetes patients in which a policy monitors glucose levels and administers insulin. The true reward captures patient health while the proxy reward prioritizes reducing the monetary cost of insulin.
@@
We also study RLHF, in which LLMs are optimized based on a reward function learned from human preferences. Following Gao et al. (2022) and Coste et al. (2024), we use a large (more robust) reward model as the true reward function and a smaller (less robust) one as the proxy. Finally, as an interpretable and illustrative example, we include the tomato-watering gridworld from Leike et al. (2017). In this environment, a robot that waters tomatoes receives true reward depending on how many plants are watered; however, the proxy additionally rewards the robot for standing in a sprinkler where it appears all the tomatoes are watered. Besides the gridworld, all of our environments reflect complex, realistic tasks with large or infinite state state spaces. To test our definition of reward hacking, we construct a natural base policy for each of the five environments. See Appendix C for more details about the environments and base policies.
@@
The top row of Figure 2 shows the distribution of true and proxy reward values for state-action pairs sampled from these base policies in each environment. We find that in all cases, the proxy rewards correlate with the true reward, satisfying our definition of a correlated proxy. The middle row of Figure 2 shows the result of optimizing for the proxies: for each environment, we plot the distribution of true and proxy reward values for state-action pairs reached by a policy optimized on the proxy reward via RL. In all cases, we see that the true reward drops significantly compared to the base policy. Thus, these intuitive cases of reward hacking are captured by Definition 4.2.
@@
Figure 3: Unlike RLHF, which attempts to prevent reward hacking by regularizing action distribution (AD) divergence from the base policy, our results suggest regularizing using occupancy measure (OM) divergence is more effective. These plots of the glucose monitoring environment show the typical ADs and OMs of two policies. $\pi$ is close to $\pi_{\text{base}}$} in AD; it is more likely to give slightly less insulin due to the proxy reward’s monetary cost penalty. However, $\pi$’s optimization of the proxy also leads to a vastly different OM with typical glucose levels far outside the healthy range (dotted lines). Thus, regularizing ADs to be close to $\pi_{\text{base}}$} is not enough to prevent reward hacking. Instead, divergence between the OMs better captures the reward hacking behavior.

### v4

@@ Abstract
Because it is difficult to precisely specify complex objectives, reinforcement learning policies are often optimized using proxy reward functions that only approximate the true goal. However, optimizing proxy rewards frequently leads to reward hacking: the optimized reward function ceases to be a good proxy and the resulting policy performs poorly with respect to the unspecified true reward. Principled solutions to reward hacking have been impeded by the lack of a good definition for the problem. To address this gap, we introduce a definition of reward hacking based on the correlation between proxy and true rewards for states and actions seen by a “reference policy” that breaks down under optimization. We show that this definition captures reward hacking behavior across several realistic settings, including in reinforcement learning from human feedback (RLHF). Using our formulation, we show theoretically that regularization to the reference policy can effectively prevent reward hacking. While the current practice in RLHF applies a KL penalty between action distributions for this purpose, our theory suggests regularizing the $\chi^{2}$ divergence between the policies’ occupancy measures can be more effective. We intuitively show the benefits of this type of regularization and demonstrate that it better mitigates reward hacking in practice across four realistic settings, including RLHF. Our code is available at https://github.com/cassidylaidlaw/orpo.
@@ 1 Introduction
Since proxy reward functions for complex tasks are nearly always misspecified in practice, what can be done to avoid reward hacking? There is a lack of principled solutions for preventing reward hacking, stemming from a more fundamental problem: defining reward hacking in a formal sense that captures realistic cases. In particular, it is difficult to characterize what makes a proxy reward function “good”. We argue that proxies are chosen because they seem to capture the true objective under some reference distribution of behavior. To formalize this, we define a proxy reward as one that correlates with the true reward function under the distribution of states and actions visited by a reference policy. Then, we define reward hacking in this setting as when optimizing a correlated proxy reward leads to a policy obtaining lower true reward than the reference policy (Figure 1). This captures several intuitive cases of reward hacking across realistic environments: traffic control, pandemic mitigation, blood glucose regulation, and reinforcement learning from human feedback (RLHF) (Figure 2).
Figure 1: We present a new characterization of reward hacking and a method for preventing it. We define a proxy reward function as one that correlates with an unknown true reward function for state-action pairs sampled from some reference policy. However, optimizing the proxy alone can lead to a breakdown in the correlation and worse true reward than the reference policy. We show theoretically and empirically that optimizing the proxy with $\chi^{2}$ occupancy measure regularization to the reference policy can allow outperforming the reference policy under the unknown true reward.
@@
Furthermore, our definition motivates a method for avoiding reward hacking during policy optimization by regularizing the optimized policy to be similar to the reference policy. Specifically, we find that optimizing the proxy reward minus a regularization term provides a provable lower bound on improvement in true reward. The amount of regularization needed increases as the correlation between the proxy and true rewards becomes weaker. In practice, RLHF already uses this form of regularization through a KL divergence penalty to the reference policy (Stiennon et al., 2020; Bai et al., 2022), and Theorem 5.1 provides theoretical justification for why this works. However, our results suggest an important modification: rather than penalizing the KL divergence between the policies’ action distributions as currently done in RLHF, we should instead penalize the chi-squared ($\chi^{2}$) divergence between the policies’ occupancy measures. We compare both OM to AD regularization and $\chi^{2}$ to KL divergence, presenting intuitive reasons why $\chi^{2}$ OM divergence may be a better regularization target in Figures 3 and 4.
Based on our theoretical results, we empirically investigate the benefits of using $\chi^{2}$ occupancy measure regularization to prevent reward hacking. Following previous work (Ho & Ermon, 2016), we implement OM regularization using a discriminator network that approximates the OM divergence between policies. We evaluate our approach in multiple reward hacking benchmark environments (Pan et al., 2022). Our results demonstrate that training with occupancy measure regularization leads to better performance under the unseen true reward function in all of the environments, validating our theoretical results. In contrast, we find that it is difficult to tune AD regularization in some environments to both prevent reward hacking and allow meaningful improvement over the reference policy. Furthermore, regularization with $\chi^{2}$ divergence tends to be more robust to the choice of regularization coefficient and often achieves higher true reward than regularizing with KL divergence.
@@
    We provide a new formal definition of reward hacking that captures realistic case studies.
@@
    Using our definition, we establish that optimizing a correlated proxy reward with $\chi^{2}$ occupancy measure (OM) regularization leads to a provable improvement in the unknown true reward function.
@@ 2 Related Work
Reward hacking.  Some prior works establish theoretical models of reward hacking as a special case of Goodhart’s Law (Goodhart, 1984; Leike et al., 2018; Manheim & Garrabrant, 2019; Krakovna, 2019; Skalse et al., 2022; Ngo et al., 2023; Fluri et al., 2024; Kwa et al., 2024). Krakovna (2018) catalog many examples of reward hacking. Pan et al. (2022) categorize different types of reward misspecification and relate optimization power to reward hacking. See the end of Section 4 for a comparison of our definition of reward hacking with previous ones.
Safe reinforcement learning.  In the context of reward hacking, regularizing policies to be similar to an offline policy based on their action distribution KL divergence was proposed by Stiennon et al. (2020) and has since been widely employed in the context of optimizing LLMs using RLHF (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022). KL regularization for RLHF has been further studied by Vieillard et al. (2021), Gao et al. (2022), and Korbak et al. (2022). Nika et al. (2024) propose a type of occupancy measure regularization in RLHF but it is limited to deterministic MDPs, while we study general stochastic MDPs. Some alternative approaches to avoid reward hacking include quantilizers (Taylor, 2016), “mild” optimization (Taylor et al., 2020), and impact regularization (Turner et al., 2020). While constrained RL can prevent the misbehavior of agents that optimize flawed reward functions (Dalal et al., 2018; Chow et al., 2019; Zhang et al., 2020; Roy et al., 2022), it simply shifts the difficulty of designing a reward function to specifying a set of constraints and weights. Robust RL usually considers a misspecified transition model, but some work has explored misspecified reward functions (Derman et al., 2021; Gadot et al., 2024). Other proposals to address the reward specification problem attempt to infer the true reward function based on the given proxy reward function, environment context, and/or feedback from humans (Hadfield-Menell et al., 2017; Reddy et al., 2020; Lee et al., 2021). Gleave et al. (2021) and Skalse et al. (2024) have previously studied quantifying the similarity of reward functions.
@@
Applications of occupancy measure regularization.  Occupancy measure regularization and optimization have been used for a variety of purposes, including imitation learning (Ho & Ermon, 2016) and efficient exploration Kang et al. (2018); Hazan et al. (2019); Lee et al. (2020); Nedergaard & Cook (2023). Various types of distributional regularization are used in model-based RL since learned models may not generalize out-of-distribution (Yang et al., 2022).
@@
Refer to caption04812-100-10-50Referencepolicy(4.5, -7.6)Trafficcontrol(r=0.10)Refer to caption04812-100-10-50Rewardhacking policyTrue reward $R(s,a)$(9.1, -197.2)Refer to caption04812-100-10-50Regularizedpolicy(4.9, -4.2)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.06, -0.06)Pandemicmitigation(r=0.75)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.01, -0.14)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.05, -0.05)Refer to caption-0.2-0.10-1000-100-500(-0.04, -28)Glucosemonitoring(r=0.70)Refer to caption-0.2-0.10-1000-100-500(-0.00, -106)Refer to caption-0.2-0.10Proxy reward $\tilde{R}(s,a)$-1000-100-500(-0.01, -8)Refer to caption-3030102030(-0.0, 16.4)RLHF(r=0.40)Refer to caption-3030102030(1.4, 10.0)Refer to caption-3030102030(0.3, 17.0)Refer to caption00.20.400.050.1(0.07, 0.06)AI safetygridworld(r=0.40)Refer to caption00.20.400.050.1(0.45, 0.02)Refer to caption00.20.400.050.1(0.09, 0.09)Mean reward
Figure 2: Our definition of reward hacking successfully describes reward hacking behavior in four realistic environments and an illustrative gridworld. The top row shows the distribution of proxy and true reward values for state-action pairs sampled from a domain-appropriate reference policy for each environment; in all environments, the proxy and true rewards are correlated. However, as shown in the middle row, this correlation breaks down if the proxy is optimized via RL, and the true reward achieved is lower than that of the reference policy, which we characterize as reward hacking. In line with our theoretical results, the bottom row shows that RL with occupancy measure regularization to the reference policy can prevent reward hacking while enabling an increase in true reward.
@@ 3 Preliminaries
To summarize, we argue that a good definition of reward hacking should satisfy two desiderata:
  1. 1.
    Characterizing what makes a proxy reward “reasonable” to optimize in the first place.
  2. 2.
    Defining a threshold for true reward below which a policy is reward hacking.
After presenting our definition in the next section, we compare to other definitions using these criteria.
@@ 4 Defining Reward Hacking
Despite the difficulties in formalizing reward hacking, we argue that both of our desiderata for a definition can be met by defining reward hacking with respect to a reference policy. This approach enables both a precise definition of proxy rewards and a natural true reward threshold for determining when reward hacking is occurring.
Characterizing proxy rewards.  To find a realistic definition of a good proxy reward, consider the process by which system designers create proxies. If they are hand-specified, we argue that designers usually imagine a distribution of behavior and design a reward that captures the objective under that distribution. For example, we study a traffic control simulator (Lopez et al., 2018; Vinitsky et al., 2018; Wu et al., 2022; Pan et al., 2022) where a small number of autonomous cars help to regulate traffic among a larger population of human drivers. In this case, a designer might choose average vehicle speed as a proxy for traffic flow; this is a reasonable proxy because under the distribution of typical human driving behavior, higher speeds are associated with better traffic conditions. We can formalize this intuition by requiring that the proxy and true rewards be correlated over states and actions sampled from a reference policy:
@@
An $r$-correlated proxy reward $\tilde{R}$ with respect to a reference policy $\pi_{\text{ref}}$} is one that has a correlation of $r>0$ with the true reward for state-action pairs sampled from the reference policy:
@@
$\textstyle\qquad\qquad\quad\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\left(\frac{\tilde{R}(s,a)-J({\pi_{\text{ref}}},\tilde{R})}{\sigma_{\tilde{R}}}\right)\left(\frac{R(s,a)-J({\pi_{\text{ref}}},R)}{\sigma_{R}}\right)\right]=r,$
@@
$\textstyle\text{where}\quad\sigma_{\tilde{R}}^{2}=\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\big(\tilde{R}(s,a)-J({\pi_{\text{ref}}},\tilde{R})\big)^{2}\right]\quad\text{and}\quad\sigma_{R}^{2}=\mathbb{E}_{\mu_{\pi_{\text{ref}}}}\left[\big(R(s,a)-J({\pi_{\text{ref}}},R)\big)^{2}\right]$
are the variances of proxy and true rewards under the reference policy.
@@
To consider an application of our definition, let’s return to the example of our traffic environment, where we define the true reward function as the negative total commute time for all vehicles. Using an autonomous vehicle policy based on a common model of human driving behavior as $\pi_{\text{ref}}$}, we plot the distribution of true and proxy rewards in the top-left corner of Figure 2. The clear correlation that we observe between the two reward functions—higher average speed corresponding to lower commute times and vice versa—confirms that average speed is a correlated proxy reward according to Definition 4.1.
@@
Definition 4.1 offers key advantages over past formalisms. First, it avoids too strongly constraining proxy rewards by allowing the proxy and true reward functions to diverge arbitrarily at some state-action pairs when those pairs have low or zero occupancy measure under the reference policy. Second, it encompasses both learned proxy rewards and hand-specified ones; see Lemma A.8 for a discussion of how learned rewards are generally correlated proxies.
@@
Choosing a threshold for reward hacking.  As discussed in Section 3, defining reward hacking requires specifying a threshold of true reward below which policy behavior is considered poor enough to be classified as reward hacking. Given that we already characterize proxy rewards with respect to a reference policy, this same reference policy provides a natural baseline for evaluating a policy that optimizes the proxy: if optimizing the proxy leads to worse true reward than the reference policy achieves, then the system designer may be better off simply using the reference policy.
Definition 4.2 (Reward hacking).
@@
Suppose $\tilde{R}$ is an $r$-correlated proxy with respect to $\pi_{\text{ref}}$} (Definition 4.1). Then we say reward hacking occurs when a policy $\pi$ optimized for $\tilde{R}$ has lower true reward than the reference policy $\pi_{\text{ref}}$}, i.e., when $J(\pi,R)<J({\pi_{\text{ref}}},R)$.
If an optimal policy for $\tilde{R}$ exhibits reward hacking, then we say that $\tilde{R}$ is a hackable correlated proxy:
@@
For example, in the traffic control environment, optimizing average speed as a proxy leads to autonomous vehicles blocking highway on-ramps, a strategy that increases overall average speed by allowing highway traffic to move freely. However, the true reward becomes extremely low because commute times for the cars on the on-ramp are arbitrarily long. This policy is thus reward hacking under Definition 4.2 since the true reward is lower than that of typical human driving.
@@
Verifying our definition experimentally.  To test whether Definition 4.2 accurately captures intuitive cases of reward hacking in practice, we consider four realistic environments and an illustrative gridworld. In addition to the aforementioned traffic simulator, we consider pandemic mitigation and glucose monitoring environments, all of which were originally studied by Pan et al. (2022) as examples of reward hacking. PandemicSimulator (Kompella et al., 2020) uses a specialized SEIR (susceptible, exposed, infected, recovered) infection model to simulate the COVID-19 pandemic among a population; the policy controls the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward omits the political cost of certain decisions. SimGlucose (Man et al., 2014) is based on an FDA-approved simulator of Type 1 Diabetes patients in which a policy monitors glucose levels and administers insulin. The true reward captures patient health while the proxy reward prioritizes reducing the monetary cost of insulin.
@@
We also study RLHF, in which LLMs are optimized based on a reward function learned from human preferences. Following Gao et al. (2022) and Coste et al. (2024), we use a large (more robust) reward model as the true reward function and a smaller (less robust) one as the proxy. Finally, as an interpretable and illustrative example, we include the tomato-watering AI safety gridworld from Leike et al. (2017). Besides the gridworld, all of our environments reflect complex, realistic tasks with large or infinite state state spaces. We construct a natural reference policy for each of the five environments; see Appendix C for more details about the environments and reference policies.
@@
The top row of Figure 2 shows the distribution of true and proxy reward values for state-action pairs sampled from these reference policies in each environment. We find that in all cases, the proxy rewards correlate with the true reward, satisfying our definition of a correlated proxy. Furthermore, when optimizing the proxies, we see that the true reward drops significantly compared to the reference policy. Thus, these intuitive cases of reward hacking are captured by Definition 4.2.
Comparison to other definitions of reward hacking.Skalse et al. (2022) define reward hacking as an increase in proxy reward accompanied by a drop in true reward. This definition does not satisfy our first desideratum of requiring the proxy to be a reasonable optimization target. Also, the threshold of any decrease in the true reward being considered reward hacking seems overly restrictive; in many cases, optimizing a proxy reward will not lead to a perfect optimization of the true reward. Kwa et al. (2024) define the phenomenon of “catastrophic Goodhart” as when optimizing a proxy with low expected error under some base policy leads to “arbitrarily high [proxy] reward despite achieving no more [true reward]” than the base policy. This definition is quite similar to ours but does not capture that the true reward decreases when reward hacking occurs; furthermore, it is often impossible to obtain arbitrarily high proxy reward in practice. Finally, Fluri et al. (2024) define error-regret mismatch as when a proxy reward has low error on a training distribution compared to the true reward, but optimizing the proxy produces a suboptimal policy. This definition is also similar to ours but leaves the threshold for reward hacking ambiguous.
Furthermore, none of these works develop a general method for preventing reward hacking. In contrast, we will show that our definition leads to a natural regularization method for optimizing a proxy reward without allowing reward hacking. Also, unlike most of these works, we validate our definition and method through extensive experiments in realistic environments.
@@
Figure 3: Unlike RLHF, which attempts to prevent reward hacking by regularizing action distribution (AD) divergence from the reference policy, our results suggest regularizing using occupancy measure (OM) divergence is more effective. These plots of the glucose monitoring environment show the typical ADs and OMs of two policies. $\pi$ is close to $\pi_{\text{ref}}$} in AD; it gives slightly less insulin. However, $\pi$’s optimization of the proxy also leads to a vastly different OM with typical glucose levels far outside the healthy range (dotted lines). Thus, regularizing ADs to be close to $\pi_{\text{ref}}$} is not enough to prevent reward hacking. Instead, divergence between the OMs better captures the reward hacking behavior.